# section[1.5em]1.5em1pc

### v2

@@ 5.   Beck modules over $A$
Let us now show that relation ($\beta$AM9) is satisfied. Let $r_{1}+\dots+r_{s}=n$, $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, let $q_{1}+\dots+q_{u}=k$, $y\in\mathcal{P}(k)^{\Sigma_{{\underline{q}}}}$. Let $(a_{i})_{i\in[s+u]}\in A^{\times s+u}$ and $m\in M$. Denote by $z=x\left(1_{\mathcal{P}}^{\times n-r_{s}},y^{\times r_{s}}\right)\in\mathcal{P}(n+r_{s}(k-1))$. On the one hand, using (**∗ ∗5), observe that:
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)\\ =\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}(\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u-1}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m))\right)\\ =\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u})),0\right)+\\ \sum_{\lambda,\lambda^{\prime}}\left(0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\left(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\right)\right),$
@@
$\sum_{\lambda,\lambda^{\prime}}\left(0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\left(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\right)\right)\\ =\sum_{\lambda,\lambda^{\prime}}\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\left(0,\beta_{\sum_{\sigma}\sigma z,({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond((1)^{\times s-1},{\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime})))}(a_{1},\dots,a_{s+u},a_{s},\dots,a_{s+u},m)\right),$
@@
$\sum_{\sigma^{\prime\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond((1)^{\times s-1},{\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime})))}}\sigma^{\prime\prime}\sum_{\sigma\in\Sigma_{({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond_{s}({\underline{q}}\circ_{u}(l,l^{\prime})))}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{(\lambda,\lambda^{\prime})}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma z\\ =\sum_{\sigma^{\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma^{\prime}z,$
@@
$\sum_{\lambda,\lambda^{\prime}}\left(0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\left(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\right)\right)\\ =\sum_{\lambda+\lambda^{\prime}=r_{s},\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}z,R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{s+u},m).$
@@
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)=$
$\left(\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})}(a_{1},\dots,a_{s+u}),0\right)+\sum_{\lambda+\lambda^{\prime}=r_{s},\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}z,R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{s+u},m).$ (5.eq.3)
@@
where $\tau$ ranges over $\Sigma_{{\underline{r}}\diamond_{s}{\underline{q}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{r_{s}}\wr\Sigma_{{\underline{q}}}$ in the sum. Using (**∗ ∗5), we then get:
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)=$
$\left(\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})}(a_{1},\dots,a_{s+u}),0\right)+\sum_{t+t^{\prime}=r_{s}q_{u},t^{\prime}>0}\left(0,\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})\circ_{s+u}(t,t^{\prime})}(a_{1},\dots,a_{s+u},m)\right).$ (5.eq.4)
Comparing the equalities (5.eq.3) and (5.eq.4), and projecting in $M$, yields the relation ($\beta$AM9).
Assigning, to each Beck module ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${A}$, the abelian $\Gamma(\mathcal{P})$-algebra $\ker(p)$ with the above $A$ module structure provides a functor:
@@
Let $M$ be an $A$-module. Consider the vector space $B:=A\oplus M$, equipped with:
@@
    For all ${\underline{r}}\in\operatorname{Comp}_{s}(n)$, $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, an operation $\beta_{x,{\underline{r}}}$ defined as in (**∗ ∗5),
@@
Assuming (A), conditions (B), (C) and (D) are obvious. To prove (A) we have to check that the operations defined in (1) satisfy the conditions ($\beta$1) to ($\beta$8). Using the notation 4.6, we can rewrite (**∗ ∗5) as:
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\dots,(a_{s},m_{s})\right)$
$\qquad=\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)+\sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{s})\right).$
@@
which, re-indexing $k=l^{\prime}$, yields
$\left(\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n)}(a),0\right)+\sum_{k=1}^{n}\left(0,\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n-k,k)}(a,m)\right)$
@@
From what precedes, this set is equal to $\coprod_{j\in[s]}\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=n-k+1\}$, which is equal to
$\coprod_{j\in[s]}\{\rho\sigma_{j}^{*}:\rho\in Sh(r_{1},\dots,r_{j-1},r_{j+1},\dots,r_{s},r_{j}-k)\times\Sigma_{1}^{\times k}\}.$
The set $Sh(r_{1},\dots,r_{j-1},r_{j+1},\dots,r_{s},r_{j}-k)\times\Sigma_{1}^{\times k}$ is empty if $k>r_{j}$, and otherwise, it is a set of representative for $\Sigma_{n-k}\times\Sigma_{k}/\Sigma_{{\underline{r}}^{\sigma_{j}}\circ_{s}(r_{j}-k,k)}$. So, we get:
$\beta_{x,{\underline{r}}}\left((a,m),\dots,(a,m)\right)=\left(\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n)}(a),0\right)+\\ \sum_{j=1}^{s}\sum_{k=1}^{n}\left(0,\beta_{\sum_{\rho\in\Sigma_{n-k}\times\Sigma_{k}/\Sigma_{{\underline{r}}^{\sigma_{j}}\circ_{s}(r_{j}-k,k)}}\rho\sigma_{j}^{*}x,(n-k,k)}(a,m)\right).$
@@
$\beta_{x,{\underline{r}}}\left((a,m),\dots,(a,m)\right)=\left(\beta_{x,{\underline{r}}}(a,\dots,a),0\right)+\\ \sum_{j=1}^{s}\sum_{k=1}^{n}\left(0,\beta_{\sigma_{j}^{*}x,{\underline{r}}^{\sigma}\circ_{s}(r_{j}-k,s)}(a,\dots,a,m)\right),$
and, using the assignment *5, we finally obtain:
@@
$\beta_{x,{\underline{r}}}\left((a,m)+(b,m^{\prime}),(a_{2},m_{2})\right)$ $=\beta_{x,{\underline{r}}}\left((a+b,m+m^{\prime}),(a_{2},m_{2})\right)$
$=\left(\beta_{x,{\underline{r}}}(a+b,a_{2}),0\right)+\sum_{l+l^{\prime}=r_{1},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime})}(a+b,m+m^{\prime},a_{2})\right)$
$\quad+\sum_{l+l^{\prime}=r_{2},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{2}(l,l^{\prime})}(a+b,a_{2},m_{2})\right),$
@@
$\sum_{k_{1}+k_{2}=r_{1}}\left(\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}(a,b,a_{2}),0\right)\\ +\sum_{l+l^{\prime}=r_{1},l^{\prime}>0}\sum_{k_{1}+k_{2}=l}\left(0,\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2},l^{\prime})}(a,b,m,a_{2})+\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2},l^{\prime})}(a,b,m^{\prime},a_{2})\right)\\ +\sum_{k_{1}+k_{2}=r_{1}}\sum_{l+l^{\prime}=r_{2},l^{\prime}>0}\left(0,\beta_{x,({\underline{r}}\circ_{2}(l,l^{\prime}))\circ_{1}(k_{1},k_{2})}(a,b,a_{2},m_{2})\right).$
@@
On the other hand, note that the assignment (**∗ ∗5) gives:
$\sum_{k_{1}+k_{2}=r_{1}}\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}\big{(}(a,m),(b,m^{\prime}),(a_{2},m_{2}$ $)\big{)}=\sum_{k_{1}+k_{2}=r_{1}}(\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}(a,b,a_{2}),0)$
$+\sum_{k_{1}+k_{2}=r_{1}}\sum_{l+l^{\prime}=k_{1},l^{\prime}>0}\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime},k_{2})}(a,m,b,a_{2}),0)$
@@
The sum of the second and third summand can be rewritten:
@@
Let us now prove relation ($\beta$8). For clarity, we can assume that $s=1$. We want to prove that:
@@
where $\tau$ ranges over $\Sigma_{(n)\diamond({\underline{q}})}/\Sigma_{(n)}\wr\Sigma_{{\underline{q}}}$ in the sum. Here, $x\in\mathcal{P}(n)^{\Sigma_{n}}$, ${\underline{q}}\in\operatorname{Comp}_{u}(k)$, and $y\in\mathcal{P}(k)^{\Sigma_{{\underline{q}}}}$. On the one hand, using the assignment (**∗ ∗5),
$\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}((a_{1},m_{1}),\dots,(a_{u},m_{u}))\right)$
$=\beta_{x,(n)}\left(\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),0\right)+\sum_{j=1}^{n}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\left(0,\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right)\right)$
$=\left(\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u})\right),0\right)+$
$\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\left(0,\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\sum_{j=1}^{u}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right)\right).$
Using relation ($\beta$8) on $A$ and ($\beta$M5) on $M$, we then get,
$\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}((a_{1},m_{1}),\dots,(a_{u},m_{u}))\right)=\left(\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}(a_{1},\dots,a_{u}),0\right)+$
$\sum_{j=1}^{u}\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\left(0,\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right)\right)$ (5.eq.5)
On the other hand, using the assignment (**∗ ∗5),
$\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}((a_{1},m_{1}),\dots,(a_{u},m_{u}))=\left(\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}(a_{1},\dots,a_{u}),0\right)+$
$\sum_{j=1}^{u}\sum_{t+t^{\prime}=nq_{j},t^{\prime}>0}\left(0,\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right).$ (5.eq.6)

### v3

@@ 5.   Beck modules over $A$
Let us now show that relation ($\beta$AM9) is satisfied. Let $r_{1}+\dots+r_{s}=n$, $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, let $q_{1}+\dots+q_{u}=k$, $y\in\mathcal{P}(k)^{\Sigma_{{\underline{q}}}}$. Let $(a_{i})_{i\in[s+u]}\in A^{\times s+u}$ and $m\in M$. Denote by $z=x\left(1_{\mathcal{P}}^{\times n-r_{s}},y^{\times r_{s}}\right)\in\mathcal{P}(n+r_{s}(k-1))$. On the one hand, using (5.eq.1), observe that:
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)$
$=\beta_{x,{\underline{r}}}\Bigg{(}(a_{1},0),\dots,(a_{s-1},0),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}(\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u-1}),\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m))\Bigg{)}$
$=\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u})),0\right)+$
$\sum_{\lambda,\lambda^{\prime}}\Bigg{(}0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\Bigg{(}a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\Bigg{)}\Bigg{)},$
@@
$\sum_{\lambda,\lambda^{\prime}}\left(0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\left(a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\right)\right)$ $=\sum_{\lambda,\lambda^{\prime}}\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\left(0,\beta_{\sum_{\sigma}\sigma z,({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond((1)^{\times s-1},{\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime})))}(a_{1},\dots,a_{s+u},a_{s},\dots,a_{s+u},m)\right),$
@@
$\sum_{\sigma^{\prime\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\Sigma_{({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond((1)^{\times s-1},{\underline{q}},{\underline{q}}\circ_{u}(l,l^{\prime})))}}\sum_{\sigma\in\Sigma_{({\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})\diamond_{s}({\underline{q}}\circ_{u}(l,l^{\prime})))}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{(\lambda,\lambda^{\prime})}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma^{\prime\prime}\sigma z\\ =\sum_{\sigma^{\prime}\in\Sigma_{R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{\lambda}\wr\Sigma_{\underline{q}}\times\Sigma_{\lambda^{\prime}}\wr\Sigma_{{\underline{q}}\circ_{u}(l,l^{\prime})}}\sigma^{\prime}z,$
@@
$\sum_{\lambda,\lambda^{\prime}}\Bigg{(}0,\beta_{x,{\underline{r}}\circ_{s}(\lambda,\lambda^{\prime})}\Bigg{(}a_{1},\dots,a_{s-1},\beta_{y,{\underline{q}}}(a_{s},\dots,a_{s+u}),\sum_{l+l^{\prime}=q_{s},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{u}(l,l^{\prime})}(a_{s},\dots,a_{s+u},m)\Bigg{)}\Bigg{)}\\ =\sum_{\lambda+\lambda^{\prime}=r_{s},\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}z,R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{s+u},m).$
@@
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)=\\ \left(\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})}(a_{1},\dots,a_{s+u}),0\right)+\sum_{\lambda+\lambda^{\prime}=r_{s},\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{u},l^{\prime}>0}\beta_{\sum_{\sigma^{\prime}}\sigma^{\prime}z,R_{\lambda\lambda^{\prime}}^{ll^{\prime}}}(a_{1},\dots,a_{s+u},m).$ (5.eq.4)
@@
where $\tau$ ranges over $\Sigma_{{\underline{r}}\diamond_{s}{\underline{q}}}/\left(\prod_{i=1}^{s-1}\Sigma_{r_{i}}\right)\times\Sigma_{r_{s}}\wr\Sigma_{{\underline{q}}}$ in the sum. Using (5.eq.1), we then get:
$\beta_{x,{\underline{r}}}\left((a_{1},0),\dots,(a_{s-1},0),\beta_{y,{\underline{q}}}((a_{s},0),\dots,(a_{s+u-1},0),(a_{s+u},m))\right)=\\ \left(\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})}(a_{1},\dots,a_{s+u}),0\right)+\sum_{t+t^{\prime}=r_{s}q_{u},t^{\prime}>0}\left(0,\beta_{\sum_{\tau}\tau z,{\underline{r}}\diamond_{s}({\underline{q}})\circ_{s+u}(t,t^{\prime})}(a_{1},\dots,a_{s+u},m)\right).$ (5.eq.5)
Comparing the equalities (5.eq.4) and (5.eq.5), and projecting onto $M$, yields the relation ($\beta$AM9).
Assigning to each Beck module ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${A}$, the abelian $\Gamma(\mathcal{P})$-algebra $\ker(p)$ with the above $A$ module structure provides a functor:
@@
Let $M$ be an $A$-module. Consider the vector space $B\coloneqq A\oplus M$, equipped with:
@@
    For all ${\underline{r}}\in\operatorname{Comp}_{s}(n)$, $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, an operation $\beta_{x,{\underline{r}}}$ defined as in (5.eq.1),
@@
Assuming (A), conditions (B), (C) and (D) are obvious. To prove (A) we have to check that the operations defined in (1) satisfy the conditions ($\beta$1) to ($\beta$8). Using Notation 4.6, we can rewrite (5.eq.1) as:
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\dots,(a_{s},m_{s})\right)=\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)+\\ \sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{s})\right).$
@@
which, re-indexing $k=l^{\prime}$, is equal to
$\left(\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n)}(a),0\right)+\sum_{k=1}^{n}\left(0,\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n-k,k)}(a,m)\right).$
@@
From what precedes, this set is equal to $\coprod_{j\in[s]}\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=n-k+1\}$, which is equal to $\coprod_{j\in[s]}\{\rho\sigma_{j}^{*}:\rho\in Sh(r_{1},\dots,r_{j-1},r_{j+1},\dots,r_{s},r_{j}-k)\times\Sigma_{1}^{\times k}\}$.
The set $Sh(r_{1},\dots,r_{j-1},r_{j+1},\dots,r_{s},r_{j}-k)\times\Sigma_{1}^{\times k}$ is empty if $k>r_{j}$, and otherwise, it is a set of representative for $\Sigma_{n-k}\times\Sigma_{k}/\Sigma_{{\underline{r}}^{\sigma_{j}}\circ_{s}(r_{j}-k,k)}$. So, the above sum is equal to:
$\beta_{x,{\underline{r}}}\left((a,m),\dots,(a,m)\right)=\\ \left(\beta_{\sum_{\sigma\in\Sigma_{n}/\Sigma_{\underline{r}}}\sigma x,(n)}(a),0\right)+\sum_{j=1}^{s}\sum_{k=1}^{n}\left(0,\beta_{\sum_{\rho\in\Sigma_{n-k}\times\Sigma_{k}/\Sigma_{{\underline{r}}^{\sigma_{j}}\circ_{s}(r_{j}-k,k)}}\rho\sigma_{j}^{*}x,(n-k,k)}(a,m)\right).$
@@
$\beta_{x,{\underline{r}}}\left((a,m),\dots,(a,m)\right)=\left(\beta_{x,{\underline{r}}}(a,\dots,a),0\right)+\sum_{j=1}^{s}\sum_{k=1}^{n}\left(0,\beta_{\sigma_{j}^{*}x,{\underline{r}}^{\sigma}\circ_{s}(r_{j}-k,s)}(a,\dots,a,m)\right),$
and, using the assignment (*5), we finally obtain:
@@
$\beta_{x,{\underline{r}}}\left((a,m)+(b,m^{\prime}),(a_{2},m_{2})\right)=\left(\beta_{x,{\underline{r}}}(a+b,a_{2}),0\right)+\\ \sum_{l+l^{\prime}=r_{1},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime})}(a+b,m+m^{\prime},a_{2})\right)+\sum_{l+l^{\prime}=r_{2},l^{\prime}>0}\left(0,\beta_{x,{\underline{r}}\circ_{2}(l,l^{\prime})}(a+b,a_{2},m_{2})\right),$
@@
$\sum_{k_{1}+k_{2}=r_{1}}\left(\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}(a,b,a_{2}),0\right)$ $+$ $\sum_{l+l^{\prime}=r_{1},l^{\prime}>0}\sum_{k_{1}+k_{2}=l}\left(0,\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2},l^{\prime})}(a,b,m,a_{2})+\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2},l^{\prime})}(a,b,m^{\prime},a_{2})\right)$ $+$ $\sum_{k_{1}+k_{2}=r_{1}}\sum_{l+l^{\prime}=r_{2},l^{\prime}>0}\left(0,\beta_{x,({\underline{r}}\circ_{2}(l,l^{\prime}))\circ_{1}(k_{1},k_{2})}(a,b,a_{2},m_{2})\right).$
@@
On the other hand, note that the assignment (5.eq.1) gives:
$\sum_{k_{1}+k_{2}=r_{1}}\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}\big{(}(a,m),(b,m^{\prime}),(a_{2},m_{2})\big{)}$
$=$ $\sum_{k_{1}+k_{2}=r_{1}}(\beta_{x,{\underline{r}}\circ_{1}(k_{1},k_{2})}(a,b,a_{2}),0)+\sum_{k_{1}+k_{2}=r_{1}}\sum_{l+l^{\prime}=k_{1},l^{\prime}>0}\beta_{x,{\underline{r}}\circ_{1}(l,l^{\prime},k_{2})}(a,m,b,a_{2}),0)$
@@
The sum of the second and third summands can be rewritten:
@@
Let us now prove relation ($\beta$8). For simplicity, we assume that $s=1$. We want to prove that:
@@
where $\tau$ ranges over $\Sigma_{(n)\diamond({\underline{q}})}/\Sigma_{(n)}\wr\Sigma_{{\underline{q}}}$ in the sum. Here, $x\in\mathcal{P}(n)^{\Sigma_{n}}$, ${\underline{q}}\in\operatorname{Comp}_{u}(k)$, and $y\in\mathcal{P}(k)^{\Sigma_{{\underline{q}}}}$. On the one hand, using the assignment (5.eq.1),
$\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}((a_{1},m_{1}),\dots,(a_{u},m_{u}))\right)=\left(\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u})\right),0\right)+\\ \sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\left(0,\beta_{x,(\lambda+\lambda^{\prime})}\left(\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\sum_{j=1}^{u}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right)\right).$
Using relation ($\beta$8) on $A$ and ($\beta$M5) on $M$, we then get:
$\beta_{x,(n)}\left(\beta_{y,{\underline{q}}}((a_{1},m_{1}),\dots,(a_{u},m_{u}))\right)=\left(\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}(a_{1},\dots,a_{u}),0\right)+\\ \sum_{j=1}^{u}\sum_{\lambda+\lambda^{\prime}=n,\lambda^{\prime}>0}\sum_{l+l^{\prime}=q_{j},l^{\prime}>0}\bigg{(}0,\beta_{x,(\lambda+\lambda^{\prime})}\bigg{(}\beta_{y,{\underline{q}}}(a_{1},\dots,a_{u}),\beta_{y,{\underline{q}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\bigg{)}\bigg{)}$ (5.eq.6)
On the other hand, using the assignment (5.eq.1),
$\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}((a_{1},m_{1}),\dots,(a_{u},m_{u}))=\left(\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),(n)\diamond({\underline{q}})}(a_{1},\dots,a_{u}),0\right)+\\ \sum_{j=1}^{u}\sum_{t+t^{\prime}=nq_{j},t^{\prime}>0}\left(0,\beta_{\sum_{\tau}\tau\cdot x\left(y^{\times n}\right),\left((n)\diamond({\underline{q}})\right)\circ_{j}(t,t^{\prime})}(a_{1},\dots,a_{j},m_{j},a_{j+1},\dots,a_{u})\right).$ (5.eq.7)